OpenStudy (anonymous):

For the function given, state the period : F(t)= 6 sin (3t- pi/6) -1 A. pi/3 B. 2pi^2/3 C. 2pi/3 D.-1

OpenStudy (anonymous):

I think its A

OpenStudy (nincompoop):

show your solution to show that it is A

OpenStudy (anonymous):

Or C.

OpenStudy (danjs):

what is a normal period of just F(t)= sin(t) ?

OpenStudy (anonymous):

A period of a periodic function is the smallest \(p\) such that: \[ f(x+p) = f(x) \]Typically we can make this simpler by letting \(x=0\): \[ f(p) = f(0) \]There are various properties of periods that we can use. For example, we know that for a simple function \(\sin(x)\) will have a period of \(2\pi\).

OpenStudy (anonymous):

t?

OpenStudy (anonymous):

And how do we know that the period will be 2pi? @wio

OpenStudy (anonymous):

Because \(2\pi\) is a full revolution, and \(\sin\) ignores revolutions.

OpenStudy (anonymous):

ok. well how do I figure this out?

OpenStudy (anonymous):

Okay, first consider the inner function inside of \(\sin\) here. It is \(3t-\pi/6\).

OpenStudy (anonymous):

We want to find \(p\), such that: \[ (3t-\pi/6)+2\pi =3(t+p)-\pi/6 \]

OpenStudy (anonymous):

To make it more clear, we have \[ g(t)=3t-\pi/6 \]And want to find the period of: \[ 6\sin(g(t))-1 \]So we want: \[ g(t+p) =g(t)+2\pi \]

OpenStudy (anonymous):

Hmm, a better way to do this is to understand that if \(f(x)\) has a period of \(p\), then \(f(ax+b)\) has a period of \(p/a\).

OpenStudy (anonymous):

The reason for this is because when \(x=p/a\), we have \[f(ax+b)=f(a(p/a)+b) = f(p+b)=f(b)\]

OpenStudy (anonymous):

Isn't the period 2pi/3? The period of a sine function is always 2pi/B, and in this case B = 3, so the period would be 2pi/3.